OpenStudy (anonymous):

For what values of x does the graph of f(x) = x − 2 sin xhave a horizontal tangent? (Use n as your integer variable. Enter your answers as a comma-separated list.)

OpenStudy (anonymous):

f'(x)=1-2cosx Thus to have a horizontal tangent, 1-2cosx=0

OpenStudy (anonymous):

x=1/3 (-π+6 n π) x=1/3 (+π+6 n π) where n s an integer

OpenStudy (anonymous):

*is